Grade 12th passAlgebra

The mean age of a group of 100 persons was found to be 32.02, later it was discovered that the age 57 was
misread as 27. Find the correct mean.

Wajeeha Kashaf
5 Years agoGrade 12th pass

1 Answer

Aditya Gupta
5 Years ago
incorrect sum of ages = 32.02*100= 3202
27 + sum of ages of rest 99 students= 3202
sum of ages of rest 99 students= 3175
correct mean= (sum of ages of rest 99 students + 57)/100= 32.32
KINDLY APPROVE :))